what is the answer to t=12 divided by 7/4ths ?

Found 2 solutions by jim_thompson5910, rfer:
Answer by jim_thompson5910(35256)

t = 12/(7/4) = 12*(4/7) = 48/7


So t = 48/7
-------------------------------------------------------------------------------------------------
If you need more help, email me at jim_thompson5910@hotmail.com

Also, please consider visiting my website: http://www.freewebs.com/jimthompson5910/home.html and making a donation. Thank you

Jim

-------------------------------------------------------------------------------------------------

Answer by rfer(16322)

t=12/(7/4)
t=12*(4/7)
t=48/7
or
t=6 6/7